# Learning Quadrotor Control From Visual Features Using Differentiable Simulation

### v1

@@ Abstract
The sample inefficiency of reinforcement learning (RL) remains a significant challenge in robotics. RL requires large-scale simulation and, still, can cause long training times, slowing down research and innovation. This issue is particularly pronounced in vision-based control tasks where reliable state estimates are not accessible. Differentiable simulation offers an alternative by enabling gradient back-propagation through the dynamics model, providing low-variance analytical policy gradients and, hence, higher sample efficiency. However, its usage for real-world robotic tasks has yet been limited. This work demonstrates the great potential of differentiable simulation for learning quadrotor control. We show that training in differentiable simulation significantly outperforms model-free RL in terms of both sample efficiency and training time, allowing a policy to learn to recover a quadrotor in seconds when providing vehicle state and in minutes when relying solely on visual features. The key to our success is two-fold. First, the use of a simple surrogate model for gradient computation greatly accelerates training without sacrificing control performance. Second, combining state representation learning with policy learning enhances convergence speed in tasks where only visual features are observable. These findings highlight the potential of differentiable simulation for real-world robotics and offer a compelling alternative to conventional RL approaches.
@@ I Introduction
Designing effective control systems for robotic tasks remains a fundamental challenge in robotics, especially in vision-based control tasks where robots are required to operate with partial observations. Traditionally, control systems have relied on conventional cascaded controllers that require explicit and precise state estimation to function properly. Alternatively, learning-based end-to-end neural network policies have been proposed, mapping observations directly to control commands without explicit state estimation. While both paradigms have demonstrated impressive performance across various applications and real-world tasks [1, 2, 3, 4, 5, 6], they also suffer from significant limitations [7].
@@
Conventional control algorithms depend heavily on accurate state estimation and can become unreliable when these estimates are noisy or inaccurate. On the other hand, learning-based methods, such as model-free reinforcement learning (RL), often exhibit high sample complexity and unstable training processes. As a result, RL approaches typically require large-scale simulations and extensive data collection, even for relatively simple state-based control tasks. Moreover, in many real-world robotic scenarios where the state is not directly observable, the sample complexity increases further as the neural network must simultaneously learn both state representations from feature observations and the control policy itself.
Despite significant efforts to enhance state estimation methods, there has been relatively less focus on developing control algorithms capable of effectively handling both state-based control tasks—where the robot’s state is known—and vision-based control tasks that rely solely on feature observations. While many control algorithms are effective in state-based control, they often struggle with controlling a robot directly without state estimation. This gap motivates our work and leads us to address the following research question: How can we design a control system that effectively solves robot control tasks in both state-based and vision-based scenarios?
We investigate this question by focusing on quadrotor control, where the objective is to control a quadrotor using only visual features extracted from camera observations. This is a particularly difficult task for several reasons. First, quadrotors are inherently underactuated and unstable systems; without continuous and precise feedback control, they quickly lose stability and crash. Second, visual features only provide information about pixel values and their positions within the image plane. These features cannot be directly applied in classical control algorithms, which typically require full-state information such as position, orientation, and velocity for solving a control task. Finally, the system must simultaneously infer both the state representation in an implicit manner and the control law from visual inputs alone, significantly complicating the optimization of the control task.
We explore differentiable simulation, an emerging learning-based strategy, for quadrotor control. Differentiable simulation provides low-variance, first-order gradients by backpropagating gradients through the dynamics model. This capability opens up significant potential for effectively optimizing neural network control policies, especially for more complex, vision-based applications and beyond.
@@
We begin with a classical state-based control task and then transition to vision-based control directly from visual features (i.e., without state estimation). Remarkably, after just 23 seconds of training, our state-based control policy trained in differentiable simulation successfully learns to stabilize the quadrotor from randomly initial configurations. In comparison, PPO (Proximal Policy Optimization) [8] requires over five times more samples and training time to achieve similar results. For feature-based control, our network trained with differentiable simulation learns to stabilize the quadrotor based on observations from visual features only. In contrast, PPO, even after extensive training, converges to a sub-optimal reward that is significantly lower than that of differentiable simulation, highlighting its limitations in vision-based tasks.
@@
Contribution: The key contribution of our work is leveraging differentiable simulation for quadrotor control, focusing on controlling the vehicle from visual features only. We demonstrate that differentiable simulation offers a significant advantage over reinforcement learning for both state-based and vision-based control tasks. Importantly, we provide detailed descriptions of our design choices and several important insights that are useful for researchers to tackle more challenging tasks. First, incorporating a simplified surrogate model in backward pass speeds up gradient computation without compromising on sample efficiency and performance. Second, pretraining the control policy on an auxiliary representation task accelerates and stabilizes the overall training process. This paper may also serve as a tutorial for researchers to apply differentiable simulation for other robot control tasks.
@@ II Related Work
One of the most common control strategies for quadrotors is the cascaded PID controller, where the inner loop controls the attitude and the outer loop controls the position or velocity [9, 10, 3, 11, 6]. Optimal control methods are also commonly used for quadrotor control, such as Linear Quadratic Regulator (LQR) and Model Predictive Control (MPC). In [12], the authors presented body-rate control using LQR to consider both the body rate and the single motor dynamics. MPC handles nonlinear systems, constraints, and disturbances more effectively than LQR. Additionally, MPC formulations are more flexible since they can incorporate different loss terms. For example, PAMPC [13] optimizes perception objectives for robust and reliable sensing by maximizing the visibility of a point of interest and minimizing its velocity in the image plane. The conventional control methods require explicit state estimation and are very sensitive to the accuracy of a given estimated state.
Neural network-based controllers offer a promising alternative to conventional methods due to their universal approximation capability, allowing the simultaneous learning of both state representation and control policies.
Model-free reinforcement learning (RL) has been used for optimizing neural network control policies for quadrotor control, including stabilizing a quadrotor [2, 14] and high-speed drone racing [15, 16]. However, the progress in applying RL to quadrotor control is largely driven by the enhanced computational capabilities provided by modern hardware rather than substantial breakthroughs in the underlying algorithms. Consequently, RL struggles with tasks where data collection is difficult to be accelerated through computational means, such as vision-based control. In general, vision-based control presents a significant challenge for RL because it requires the system to learn a state representation alongside the control policy from high-dimensional observations.
Differentiable simulation is an emerging technique in robot learning that enables the propagation of gradients through a system’s dynamics model, allowing for more efficient policy optimization via analytical policy gradients. Most studies on differentiable simulation for policy optimization have been limited to simulations [17, 18, 19], with only a few successes in real-world applications, such as learning quadruped locomotion [20] and agile, vision-based drone swarm navigation [21]. In particular, [21] demonstrates how a simple point-mass model can be used to learn vision-based navigation policies for agile flight in complex real-world environments. Despite its success, this approach does not account for the vehicle’s rotational dynamics. There remains a gap in research exploring low-level quadrotor control that fully incorporates the complete vehicle dynamics, which is important for leveraging the full potential of a quadrotor for super agile maneuvers, such as acrobatic flight or time-optimal flight.
@@ III-A Policy Optimization via Differentiable Simulation
The quadrotor is modeled as a discrete-time dynamical system, characterized by continuous state and control input spaces, denoted as $\mathcal{X}$ and $\mathcal{U}$, respectively. At each time step $t$, the system state is $x_{t}\in\mathcal{X}$, and the corresponding control input is $u_{t}\in\mathcal{U}$. An observation $o_{t}\in\mathcal{O}$ is generated at each time step based on the current state $x_{t}$ through an observation model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{t}=h(x_{t})$. The system dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system $x_{t+1}=f(x_{t},u_{t})$. At each time step $t$, the robot receives a reward signal $r_{t}=r(x_{t},u_{t})$. The control policy is a deterministic, differentiable function, such as a neural network, $u_{t}=\pi_{\theta}(o_{t})$. The neural network takes the observation $o_{t}$ as input and outputs the control input $u_{t}$.
The objective function, the cumulative reward given the policy paramters $R(\theta)$, is to be maximized to find the optimal policy parameters $\theta^{\ast}$ by means of gradient ascent
@@
Back-propagation Through Time (BPTT): Based on the structure of $R(\theta)$, the policy gradient is obtained through BPTT and reads
@@
$\frac{1}{N}$ $\sum_{t=0}^{N-1}\left(\sum_{i=1}^{t}\frac{\partial r_{t}}{\partial x_{t}}\prod_{j=i}^{t}\left(\frac{dx_{j}}{dx_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial r_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$
where the matrix of derivatives $dx_{j}/dx_{j-1}$ is the Jacobian of the close-loop dynamical system $f(x,\pi_{\theta}(h(x)))$.
@@ III-B Quadrotor Dynamics
Let $\bm{p}$, $\bm{R}$, and $\bm{v}$ represent the position, orientation matrix, and linear velocity of the quadrotor, respectively, expressed in the world frame. Let $\bm{\omega}$ denote the angular velocity of the quadrotor expressed in the body frame. Additionally, let $c$ represent the mass-normalized collective thrust produced by all motors, and let $\bm{c}=\begin{bmatrix}0&0&c\end{bmatrix}^{\intercal}$ denote the collective thrust vector and $\bm{g}$ is the gravity vector. Finally, $\bm{J}$ is the diagonal moment of inertia. The simple quadrotor dynamics read
$\dot{\bm{x}}=\begin{bmatrix}\bm{p}\\ \operatorname{vec}(\bm{R})\\ \bm{v}\\ \end{bmatrix}=\begin{bmatrix}\bm{v}\\ \operatorname{vec}(\mathbf{R}[\bm{\omega}]_{\times})\\ \bm{R}\bm{c}+\bm{g}\\ \end{bmatrix}$ (1)
where $[\cdot]_{\times}$ is the skew symmetric matrix operator and $\operatorname{vec}(\cdot)$ indicates the vectorization of the input matrix.
@@ III-C JAX-based Differentiable Simulator
To enable back-propagating gradients through simulator rollouts and exploit GPU-accelerated computing, the differentiable simulation framework is entirely written using JAX [22]. During the runtime, the code is just-in-time compiled and lowered to GPU, resulting in fast execution times of up to million of steps per second (Table I). Besides high learning speeds, JAX supports automatic differentiation which can be used to compute analytical policy gradients.
@@ III-D Fast Surrogate Gradients
TABLE I: Simulations speeds with and without gradient computation (BPTT) on an Nvidia Titan RTX (24 GB VRAM). Using a simple model for the backward path (BPTT simple) keeps computation and memory costs low. For empty entries, the computational graph required more memory than accessible.
@@ III-E Pretraining on State Representation Learning
The pretraining procedure comprises three steps. First, a randomly initialized policy network is used to collect data. Then, a neural network is trained to predict the underlying simulation state (the quadrotor state) from the visual observations. Lastly, the resulting weights are copied to the policy network.
@@ III-F1 Policy
We parameterize the policy as a multi-layer perceptron with two hidden layers. The size of the hidden layers is 512 state-based and 1024 for vision-based control.
@@ III-F4 Reward Function
Stabilizing the quadrotor means regulating its state toward a hovering condition. The desired behavior is encoded into the reward function. At time $t$
$r_{t}=r_{t}^{\mathrm{pos}}+r_{t}^{\mathrm{vel}}+r_{t}^{\mathrm{act}}$
@@
$r_{t}^{\mathrm{pos}}$ $=-0.2\cdot L_{\mathrm{H}}(5\cdot(\bm{p}_{t}-\bm{g}))$
@@
Here, $L_{\mathrm{H}}$ is the Huber loss and $\bm{u}_{\mathrm{hover}}=[9.81,0,0,0]^{\top}$, the action required to withstand the gravity.
@@ IV-A Learning State-based Control
The table in Figure 4 compares how many samples and how much training time is needed to reach certain reward targets. We find that BPTT reaches all targets significantly faster and requires only a fraction of samples. In particular, BPTT demonstrates an almost 8-times speed-up to reach the highest reward target -5.3 and requires less than 20% of samples.
@@ IV-B Feature-based Control without State Estimation
For the more challenging task of vision-based quadrotor control, we pretrain the networks using state representation learning for about 1:30 min and subsequently run policy optimization for 2000 iterations. Unsurprisingly, the time to convergence is higher in all cases. Only BPTT with 100 or 1000 parallel environments manages to converge to a performance close to the state-based case. Even disregarding the significantly lower reward of PPO policies at convergence, the convergence time of BPTT, 9:11 min, is more than 3 times lower compared to 32:06 min, the convergence time of PPO. Furthermore, The table in Figure 5 displays the wall-clock training time and simulation samples required to reach the reward targets. Again, we observe much lower sample and time requirements. For the last target hit by PPO, a reward of -11, BPTT was over 70 times faster and needed less than 4% of samples.

### v2

@@ Abstract
The sample inefficiency of reinforcement learning (RL) remains a significant challenge in robotics. RL requires large-scale simulation and can still cause long training times, slowing research and innovation. This issue is particularly pronounced in vision-based control tasks where reliable state estimates are not accessible Differentiable simulation offers an alternative by enabling gradient back-propagation through the dynamics model, providing low-variance analytical policy gradients and, hence, higher sample efficiency. However, its usage for real-world robotic tasks has yet been limited. This work demonstrates the great potential of differentiable simulation for learning quadrotor control. We show that training in differentiable simulation significantly outperforms model-free RL in terms of both sample efficiency and training time, allowing a policy to learn to recover a quadrotor in seconds when providing vehicle states and in minutes when relying solely on visual features. The key to our success is two-fold. First, the use of a simple surrogate model for gradient computation greatly accelerates training without sacrificing control performance. Second, combining state representation learning with policy learning enhances convergence speed in tasks where only visual features are observable. These findings highlight the potential of differentiable simulation for real-world robotics and offer a compelling alternative to conventional RL approaches.
@@ I Introduction
Reinforcement Learning (RL) has become a cornerstone in robotics research for designing powerful controllers. By employing large-scale simulation, neural network controllers trained with RL enabled impressive performance in many robotic tasks [1] like quadruped locomotion [2], robot soccer [3], or quadrotor control [4, 5]. The learning paradigm allows for the end-to-end design of control policies that take various input modalities and optimize task-level objectives [6]. However, the flexibility comes at the cost of low sample efficiency, which slows down the development of new controllers. Successful implementations require massively parallelized simulators and large compute resources. Yet, training times can go up to hours or even days. This issue makes iterating on the control architecture, environment settings, and hyperparameter tuning tedious and delays new results and progress.
The sample efficiency problem becomes even more pronounced in vision-based control tasks [7]. Not only is the simulation more expensive due to the additional requirement of simulating the visual observations, a direct mapping from visual input signals to control commands is more difficult to learn. First, the dimensionality of the inputs increases, requiring more samples to optimize the policy. Second, the neural network must simultaneously build an implicit state representation and synthesize the control law.
Differentiable simulation offers an approach to improve the sample efficiency in robot learning. By incorporating the derivatives of the dynamics, policy gradients can be computed analytically. This method reduces the variance in the policy parameter updates, which potentially decreases the number of samples needed to learn a new control task.
@@
We investigate using differentiable simulation to learn vision-based quadrotor control. In our case study, the objective is to stabilize a quadrotor from a manual throw using only visual features extracted from camera observations (see Figure 1). The task is challenging for several reasons. First, quadrotors are inherently underactuated and unstable systems; without continuous and precise feedback control, they quickly lose stability and crash. Second, state information is only indirectly available through pixel coordinates in the image plane, making the environment partially observable. Finally, the control policy has to handle a wide initial state distribution to recover from a manual throw.
@@
In our study, we begin with a classical state-based control task and then transition to vision-based control directly from visual features (i.e., without state estimation). We compare learning in differentiable simulation with PPO (Proximal Policy Optimization) [8], a popular model-free RL algorithm. Remarkably, after just 23 seconds of training, our state-based control policy trained in differentiable simulation successfully learns to stabilize the quadrotor from random initial configurations. In comparison, PPO requires over five times more samples and training time to achieve similar results. For feature-based control, our network trained in differentiable simulation learns to stabilize the quadrotor based on observations from visual features only. In contrast, PPO, even after extensive training, converges to a sub-optimal reward that is significantly lower than that of differentiable simulation, highlighting its limitations in vision-based tasks.
@@
Contribution: The key contribution of our work is leveraging differentiable simulation for quadrotor control, focusing on controlling the vehicle from visual features only. We demonstrate that, first, differentiable simulation offers a significant advantage over model-free RL for both state-based and vision-based control tasks. Second, incorporating a simplified surrogate model in backward pass speeds up gradient computation without compromising on sample efficiency and performance. Third, pretraining the control policy on an auxiliary representation task accelerates and stabilizes the overall training process.
@@ II Related Work
Model-free RL is popular method to learn quadrotor control policies, including stabilization [4, 9] and high-speed drone racing [6, 10]. For vision-based control, it is common to split the design process into state estimator and controller design. In [10], the quadrotor state is estimated using a Kalman filter that receives measurements from a visual-inertial odometry module and additional visual features. The state estimate is then fed into a state-based policy that was trained using RL. Splitting state estimation and control gives rise to a modular structure. However, the composition can lead to new failure modes caused by compounding errors. Alternatively, an end-to-end learning approach is studied in [7] for drone racing. The authors demonstrate robust flight performance solely relying on visual inputs. Their experiments show that learning from visual inputs requires much more samples and longer training times of around 24 hours.
Most studies on differentiable simulation for policy optimization in robotics have been limited to simulated experiments [11, 12, 13], with only a few successes in real-world applications, such as learning quadruped locomotion [14] and vision-based drone swarm navigation [15]. In particular, [15] demonstrates how a simple point-mass model can be used to learn vision-based navigation policies for agile flight in complex real-world environments. Their approach requires state estimation and a low-level controller that can stabilize the quadrotor and track the acceleration commands from the policy. Hence, there remains a research gap in exploring low-level quadrotor control that fully incorporates the complete vehicle dynamics, which is important for leveraging the full potential of a quadrotor for agile maneuvers, such as acrobatic flight or time-optimal flight.
Using a surrogate model for policy gradient computations was first studied in [16]. The authors propose using an approximate simulation model to compute gradients along real-world trajectories. They prove that the resulting policy parameters lie close to a local optimum with respect to the true but unknown model. Another work [17] even considers the case in which the surrogate model only predicts the sign of the derivatives. Recently, [14] proposed a surrogate model for gradient computations in the context of quadruped locomotion.
@@ III-A Policy Optimization via Differentiable Simulation
The quadrotor is modeled as a discrete-time dynamical system, characterized by continuous state and control input spaces, denoted as $\mathcal{X}$ and $\mathcal{U}$, respectively. At each time step $t$, the system state is $x_{t}\in\mathcal{X}$, and the corresponding control input is $u_{t}\in\mathcal{U}$. An observation $o_{t}\in\mathcal{O}$ is generated at each time step based on the current state $x_{t}$ through an observation model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{t}=h(x_{t})$. The system dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system $x_{t+1}=f(x_{t},u_{t})$. At each time step $t$, the robot receives a reward signal $r_{t}=r(x_{t},u_{t})$. The control policy is a deterministic differentiable function, a neural network, $u_{t}=\pi_{\theta}(o_{t})$. The policy takes the observation $o_{t}$ as the input and outputs the control input $u_{t}$.
The objective function $R(\theta)$, the cumulative reward given the policy parameters, is to be maximized to find the optimal policy parameters $\theta^{\ast}$ by means of gradient ascent
@@
Back-propagation Through Time (BPTT): By taking advantage of the structure of $R(\theta)$ and applying the chain rule along the temporal axis, the policy gradient is obtained through BPTT and reads
@@
$\frac{1}{N}$ $\sum_{t=0}^{N-1}\left(\sum_{i=1}^{t}\frac{\partial r_{t}}{\partial x_{t}}\prod_{j=i}^{t}\left(\frac{dx_{j}}{dx_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial r_{t}}{\partial u_{t}}\frac{\partial u_{t}}{\partial\theta}\right),$
where the matrix of derivatives $dx_{j}/dx_{j-1}$ is the Jacobian of the close-loop dynamical system $f(x,\pi_{\theta}(h(x)))$. For a derivation of the BPTT gradient, the reader is referred to [18].
@@ III-B Quadrotor Dynamics
Let $\bm{p}$, $\bm{R}$, and $\bm{v}$ represent the position, orientation matrix, and linear velocity of the quadrotor, respectively, expressed in the world frame. Let $\bm{\omega}$ denote the angular velocity of the quadrotor expressed in the body frame, referred to as body rates. Additionally, let $c$ represent the mass-normalized collective thrust produced by all motors, and let $\bm{c}=\begin{bmatrix}0&0&c\end{bmatrix}^{\intercal}$ denote the collective thrust vector and $\bm{g}$ the gravity vector. The simple quadrotor dynamics read
$\dot{\bm{x}}=\frac{d}{dt}\begin{bmatrix}\bm{p}\\ \operatorname{vec}(\bm{R})\\ \bm{v}\\ \end{bmatrix}=\begin{bmatrix}\bm{v}\\ \operatorname{vec}(\mathbf{R}[\bm{\omega}]_{\times})\\ \bm{R}\bm{c}+\bm{g}\\ \end{bmatrix},$ (1)
where $[\cdot]_{\times}$ is the skew symmetric matrix operator and $\operatorname{vec}(\cdot)$ indicates vectorization. For quadrotor control, the body rates $\bm{\omega}$ and the collective thrust $c$ are the inputs to the system.
@@ III-C JAX-based Differentiable Simulator
To enable back-propagating gradients through simulated rollouts and exploit GPU-accelerated computing, the differentiable simulation framework is entirely written using JAX [19]. During runtime, the code is just-in-time compiled and lowered to GPU, resulting in fast execution times of up to million of steps per second (Table I). Besides high learning speeds, JAX supports automatic differentiation which can be used to compute analytical policy gradients.
@@ III-D Fast Surrogate Gradients
TABLE I: Simulation speeds with and without gradient computation (BPTT) on an Nvidia Titan RTX (24 GB VRAM). Using a simple model for the backward path (BPTT simple) keeps computation and memory costs low. For empty entries, the computational graph required more memory than accessible.
@@ III-E Pretraining on State Representation Learning
The pretraining procedure comprises three steps. First, a randomly initialized policy network $\pi_{\theta}$ is used to collect a dataset $\mathcal{D}$ of pairs of simulation states and observations $(x_{t},o_{t})$. Then, a neural network is trained to predict the underlying simulation state (the quadrotor state) from the visual observations. I.e., the neural network $\psi_{\theta}$ is trained to minimize the mean squared error
$\frac{1}{N}\sum_{(x_{i},o_{i})\in\mathcal{D}}\|x_{i}-\psi_{\theta}(o_{i})\|^{2}.$
Lastly, the resulting weights $\theta$ are copied to the policy network $\pi_{\theta}$. Note that the last layer is changed do to the different output modalities.
@@ III-F1 Policy
We parameterize the policy as a multi-layer perceptron with two hidden layers. The size of the hidden layers is 512 for state-based and 1024 for vision-based control.
@@ III-F4 Reward Function
Stabilizing the quadrotor means regulating its state toward a hovering condition at a desired position $\bm{p}_{\mathrm{des}}$. The desired behavior is encoded into the reward function. At time $t$
$r_{t}=r_{t}^{\mathrm{pos}}+r_{t}^{\mathrm{vel}}+r_{t}^{\mathrm{act}}.$
@@
$r_{t}^{\mathrm{pos}}$ $=-0.2\cdot L_{\mathrm{H}}(5\cdot(\bm{p}_{t}-\bm{p}_{\mathrm{des}}))$
@@
Here, $L_{\mathrm{H}}$ is the Huber loss and $\bm{u}_{\mathrm{hover}}=[9.81,0,0,0]^{\top}$, the action required to withstand the gravity. We use the Huber loss to limit the magnitude of the gradient even for large deviations to improve training stability.
@@ IV-A Learning State-based Control
The table in Figure 4 compares how many samples and how much training time is needed to reach certain reward targets (with 1000 parallel environments). We find that BPTT reaches all targets significantly faster and requires only a fraction of samples. In particular, BPTT demonstrates an almost 8-times speed-up to reach the highest reward target -5.3 and requires less than 20% of samples.
@@ IV-B Feature-based Control without State Estimation
For the more challenging task of vision-based quadrotor control, we pretrain the policy networks using state representation learning on 100,000 samples for 500 epochs (pretraining time is around 1:30 min) and subsequently run policy optimization for 2000 iterations. Unsurprisingly, the time to convergence is higher in all cases. Only BPTT with 100 or 1000 parallel environments manages to converge to a performance close to the state-based case. Even disregarding the significantly lower reward of PPO policies at convergence, the convergence time of BPTT, 9:11 min, is more than 3 times lower compared to 32:06 min, the convergence time of PPO. Furthermore, the table in Figure 5 displays the wall-clock training time and simulation samples required to reach the reward targets. Again, we observe much lower sample and time requirements. For the last target hit by PPO, a reward of -11, BPTT was over 70 times faster and needed less than 4% of samples.